An auto manufacturer claims that its new 4-cylinder Hybrid auto with manual shift averages 50 mpg under city driving conditions. A Federal Agency believes this claim is too high and decides to randomly test 25 of the manufacturer's 4-cylinder Hybrid autos with manual shift. The Agency determines the mean average mpg for the sample to be 48 with a sample standard deviation of 4 mpg under city driving conditions. The p-value for this hypothesis test is:
Answer #1

Solution :

This is the left tailed test,  

The null and alternative hypothesis is ,

H0 :  \mu = 50

Ha : \mu < 50

Test statistic = t

= (T - \mu ) / s / \sqrt n

= (48 - 50) / 4 / \sqrt 25

Test statistic = t = -2.50

degrees of freedom = n - 1 = 25 - 1 = 24

P(t < -2.50 )

P-value = 0.0098